how do I work this problem out? Solve for x: 4/x=1/5
Answer by stanbon(75887)

4/x=1/5
----------
You could cross-multiply, then solve for "x".
Or you could invert both sides to get:
x/4 = 5/1
Multiply both sides by 4 to get x = 20
=================
Cheers,
Stan H.